The knobcone gene has two alleles that segregate within a population of loblolly pine. In this population of loblolly pine, you observe two genotypes with frequencies of 0.75 and 0.25. The fitness of the genotype with a frequency of 0.75 is equal to 1.0, whereas the fitness of the other genotype is 0.8. What is the mean fitness at this locus within this population?

Answer: The answer is 0.95. I'm not sure how to get there
Answer #1

​​​​​​​​​​​​Mean fitness = summation (genotypic frequency × fitness)

Genotypic frequency 1 = 0.75 Fitness 1 = 1

Genotypic frequency 2 = 0.25 Fitness 2 = 0.8

Mean fitness = (0.75 × 1) + (0.25 × 2) = 0.75 + 0.2 = 0.95

Answer is 0.95.

Please rate high.